# Large Language Models for Software Engineering: A Systematic Literature Review

### v5

@@ 5.3. How are evaluation metrics utilized to assess the performance of LLM4SE tasks?
In generation tasks, metrics like BLEU, along with its variants BLEU-4 and BLEU-DC (Ahmed et al., 2023; Al-Kaswan et al., 2023; Arakelyan et al., 2023; Chen et al., 2022; Ciniselli et al., 2021), and Pass@k (Bui et al., 2023; Cassano et al., 2023; Chen et al., 2023c, 2021b; Dibia et al., 2022; Döderlein et al., 2022) are the most commonly used, appearing in 29 and 28 studies, respectively. For instance, Wang et al. (Wang et al., 2023c) employed BLEU to evaluate a code-to-code translation model. Pass@k is used in the research by Jiang et al. (Jiang et al., 2023a) to assess code generation models, measuring the proportion of generated code snippets that match the reference solutions. Additionally, ROUGE/ROUGE-L (Ahmed et al., 2023; Al-Kaswan et al., 2023; Gao et al., 2023; Geng et al., 2024; Li et al., 2022f; Mastropaolo et al., 2021b, 2022b; Niu et al., 2022; Zan et al., 2023a; Li et al., 2023a), METEOR (Ahmed et al., 2023; Al-Kaswan et al., 2023; Chen et al., 2022; Gao et al., 2023; Niu et al., 2022; Geng et al., 2024), EM (Exact Match) (Al-Kaswan et al., 2023; Gao et al., 2023; Gupta et al., 2023; Murali et al., 2023; Wang et al., 2023c; Weyssow et al., 2023; Ye et al., 2023; Zhang et al., 2023c), and ES (Edit Similarity) (Liu et al., 2023e) are used in specific studies to evaluate the quality and accuracy of generated code or natural language code descriptions.
RQ3 - Summary (1) We conducted an analysis of the parameters and learning rate optimizers commonly employed in LLMs, discovering that fine-tuning and Adam stand out as the most frequently utilized techniques for parameter optimization and learning rate adjustment, respectively. (2) We highlighted the application and effectiveness of prompt engineering techniques in improving the performance of LLMs for SE tasks. By exploring various types of prompts, including hard and soft prompts, this emerging fine-tuning paradigm has shown to be particularly advantageous in tasks characterized by data scarcity, providing task-relevant knowledge and enhancing LLMs’ versatility and efficacy across different code intelligence tasks. (3) We summarized the most widely used evaluation metrics according to four problem types, i.e., regression, classification, recommendation, and generation. Fifteen different evaluation metrics appeared in the generation task, while nine metrics were used for the classification task.
@@ 6.1. What are the distributions SE activities and problem types addressed to date with LLM4SE?
The highest number of studies is observed in the software development domain, constituting approximately 58.37% of the total research volume. This underscores the primary focus to date on utilizing LLMs to enhance coding and development processes. Software maintenance tasks account for about 24.89% of the research share, highlighting the significance of LLMs in aiding software updates and improvements. The software quality assurance domain holds approximately 10.3% of the research proportion, indicating a growing interest in automating testing procedures. In contrast, requirements engineering and software design activities represent approximately 4.72% and 1.29% of the research share, respectively, suggesting relatively limited exploration so far in these areas. The software management domain has the least research representation, accounting for a tiny 0.43% proportion. This distribution underscores the vital focus on development and maintenance tasks while also indicating potential avenues for further research in testing, design, and management domains.
In our collection of LLM studies for SE tasks, we’ve classified them based on the type of problems they address (shown in Fig.8 (b)). The distribution reveals that the majority of studies, about 64.34%, center around generation tasks, showcasing the significance of LLMs in producing code or text. Following this, around 24.48% of studies fall under classification tasks, indicating the relevance of LLMs in categorizing software elements. Additionally, roughly 9.79% of studies are related to recommendation tasks, demonstrating the utility of LLMs in suggesting solutions. Lastly, a smaller portion, around 1.4%, is allocated to regression tasks, reflecting the limited exploration of LLMs for predictive modeling. This distribution underscores the broad applicability of LLMs across different SE challenges, with a notable emphasis on code generation and classification tasks.
@@
Requirements engineering Anaphoric ambiguity treatment (3) Requirements term identification (1) 11
Requirements classification (3) Coreference detection (1)
Requirement analysis and evaluation (2) Traceability automation (1)
Software design GUI retrieval (1) Software specification synthesis (1) 3
Rapid prototyping (1)
Software development Code generation (62) Agile story point estimation (1) 136
Code completion (16) API documentation smell detection (1)
Code summarization (10) API entity and relation extraction (1)
Code understanding (7) Code optimization (1)
Code search (5) Code example recommendation (1)
Program synthesis (5) Control flow graph generation (1)
API recommendation (2) Data analysis (1)
API synthesis (2) Identifier normalization (1)
Code comment generation (2) Instruction generation (1)
Code representation (2) Type inference (1)
Method name generation (2) Others (11)
Software quality assurance Test generation (8) Bug localization (1) 24
Vulnerability detection (7) Failure-inducing test identification (1)
Test automation (4) Flaky test prediction (1)
Verification (2)
Software maintenance Program repair (23) Duplicate bug report detection (1) 58
Code review (6) Decompilation (1)
Debugging (4) Program merge conflicts repair (1)
Bug report analysis (3) Sentiment analysis (1)
Code clone detection (3) Tag recommendation (1)
Logging (2) Vulnerability repair (1)
Bug prediction (1) Commit classification (1)
Bug triage (1) Traceability recovery (1)
Bug report replay (1) Others (6)
Software management Effort estimation (1) 1
@@ 6.2. How are LLMs used in requirements engineering?
Anaphoric ambiguity treatment. Ambiguity in software requirements arises when a single reader can interpret a natural language (NL) requirement in multiple ways, or different readers have varying understandings of the same requirement. Unclear and ambiguous NL software requirements can lead to suboptimal software artifacts during later development stages. Moharil et al. (Moharil and Sharma, 2023) and Ezzini et al. (Ezzini et al., 2022) have empirically demonstrated the significant role of LLMs such as BERT and SpanBERT in effectively addressing anaphoric ambiguity. Sridhara et al. (Sridhara et al., 2023) revealed that ChatGPT excels in addressing anaphoric ambiguity in software requirements. Through researchers’ analysis of ten English requirement specifications (Ezzini et al., 2022) containing anaphora-related challenges, ChatGPT consistently demonstrated its remarkable capability to accurately identify antecedents. This empirical evidence emphasizes the valuable role ChatGPT can play in enhancing the clarity and precision of software requirements, ultimately contributing to more effective software development processes by reducing interpretational uncertainties.
@@ 6.4. How are LLMs used in software development?
Code generation. Code generation has long been a task of interest: there is extensive work on program synthesis using symbolic and neural-semiotic approaches (Alur et al., 2013; Wu et al., 2023b). Recently, LLMs trained for text generation have demonstrated the ability to complete programs (Brown et al., 2020; Black et al., 2022). Since 2020, several code generation models have been trained or fine-tuned on programming language text (Nijkamp et al., 2022b; Chen et al., 2021b; Fried et al., 2022; Xu et al., 2022; Feng et al., 2020; Clement et al., 2020). Unlike traditional program synthesis techniques, neurolinguistic models can be conditioned on natural language (e.g., code annotations) as well as generate programming language text. Researchers have experimentally demonstrated that LLMs like GPT-4 (Bareiß et al., 2022; Liu et al., 2023d; Jiang et al., 2023b; Gilbert et al., 2023), GPT-2/GPT-3/GPT-3.5 (Azaria et al., 2023; Yetiştiren et al., 2023; Ke et al., 2023; Liu et al., 2023d; Nascimento et al., 2023; Li et al., 2023b; Wang et al., 2023d; Liu et al., 2023a; Dong et al., 2023), BERT series (Zeng et al., 2022; Lai et al., 2023), Codex (Dibia et al., 2022; Bareiß et al., 2022; Yu et al., 2023; Chen et al., 2021b; Gupta et al., 2023; Madaan et al., 2022; Kuznia et al., 2022), CodeGen (Dibia et al., 2022; Jones and Steinhardt, 2022; Zan et al., 2022a), InCoder (Murali et al., 2023; Kou et al., 2023b; Liu et al., 2023d; Wang et al., 2022b), Copilot (Wu et al., 2023b) and CodeGeeX (Zheng et al., 2023b), play a key role in code generation. By pre-training on large-scale text data, these models learn rich linguistic knowledge and semantic representations that enable them to understand the meaning and structure of natural language. LLMs can automate code generation by converting natural language descriptions into code (Jiang et al., 2023a). These models generate program code from natural language descriptions, enhancing code-writing efficiency and accuracy. They show excellent performance in code completion, automatic code generation, and conversion of natural language annotations to code, providing software developers with powerful auxiliary tools and promoting further automation and intelligence in the code writing and development process.
@@
Within the domain of LLMs applied to software development tasks, studies centered on code generation distinctly dominate the academic landscape. As reflected in Table 12, the GPT series, particularly GPT-4, emerge as a key focus, with many more studies using them in the realm of code generation (Dong et al., 2023; Du et al., 2023; Li et al., 2023b; Liu et al., 2023d). Analysing these studies, several noteworthy findings surface:
@@
  • Programming thinking in LLMs. Techniques that evoke “programming thinking” within LLMs, such as the TIP (i.e., Thinking in Programming) (Li et al., 2023b) methodology, have shown promising strides. By guiding LLMs to first craft a high-level code sketch before delving into detailed implementations, the synthesized code exhibits higher accuracy and robustness.
@@
  • Class-level vs. Method-level generation. LLMs, while adept at method-level code generation, present varied performance metrics when tasked with class-level generation (Du et al., 2023). This divergence underscores the evolving nature of challenges as the granularity of code synthesis shifts.
@@
  • Expanding LLM capabilities. The next frontier in this discipline seems to lie in harmoniously integrating LLMs with established SE tools and practices. The emergence of frameworks like EvalPlus (Dong et al., 2023) indicates a trend towards enhancing the evaluation and accuracy of LLM-generated code, possibly ushering in an era where human developers and LLMs collaboratively craft software solutions.
@@
GPT-3.5 Codex, CodeGen, CodeGeeX, LLaMA, InCoder, PyCodeGPT, CodeParrot, GPT-2 HumanEval, MBPP, MBCPP Pass@k May 11, 2023 (Li et al., 2023b)
GPT-4 PaLM Coder, Codex, CodeGen-Mono, Incoder, CodeGeeX, AlphaCode HumanEval, HumanEval-ET, MBPP, MBPP-ET Pass@k May 24, 2023 (Dong et al., 2023)
GPT-4 GPT-3.5, StarCoder, CodeGen, CodeGen2, Vicuna, SantaCoder, Incoder, GPT-J, GPT-Neo, PolyCoder, StableLM HumanEval, HumanEval+, HumanEval-mini Pass@k Jun 12, 2023 (Liu et al., 2023d)
GPT-4 GPT-3.5, WizardCoder, Instruct-StarCoder, SantaCoder, Instruct-CodeGen, CodeGeeX, InCoder, Vicuna, ChatGLM, PolyCoder ClassEval, HumanEval Pass@k Aug 3, 2023 (Du et al., 2023)
@@
With the support of LLMs, code completion achieves significant improvements in efficiency and accuracy. Developers can save time by avoiding manual input of lengthy code and reducing the risk of code errors. LLMs also learn from extensive code repositories, acquiring knowledge and best practices to offer more intelligent and precise suggestions, aiding developers in better understanding and utilizing code (Ciniselli et al., 2021). Additionally, these models can provide personalized code recommendations based on developers’ coding styles and preferences, further enhancing the effectiveness and user experience of code completion (Liu et al., 2023e).
Code summarization. Code summarization is a task that attempts to understand the code and automatically generate descriptions directly from the source code. It can also be viewed as an extended form of documentation. Successful code summarization not only facilitates the maintenance of source code (Iyer et al., 2016; Nguyen and Nguyen, 2017) but can also be used to improve the performance of code search using natural language queries (Nie et al., 2016; Yang et al., 2016) and code classification (Nguyen and Nguyen, 2017). LLMs play a significant role in code summarization by analyzing code structures and contexts to generate informative natural language summaries. Specifically, LLMs such as Codex (Gao et al., 2023; Arakelyan et al., 2023; Ahmed et al., 2023), CodeBERT (Chen et al., 2022; Gu et al., 2022; Gao et al., 2023), and T5 (Mastropaolo et al., 2021b, 2022b) comprehend the functionality and logic of the code, producing easily understandable human language descriptions. For example, Arakelyan et al. (Arakelyan et al., 2023) rigorously evaluate the efficacy of CodeT5 and Codex across code generation and summarization tasks, shedding light on their performance under distribution shifts. It unveils practical adaptation techniques, underscoring Codex’s commendable performance. Additionally, the study demonstrates that while adapted models exhibit proficiency in code generation, their generality can present trade-offs in the context of code summarization. As a result, code summarization with the support of LLMs enhances code readability, improves software documentation quality, and accelerates code comprehension and collaboration among developers. This advanced approach to code summarization demonstrates great potential for automating and streamlining various aspects of software development in modern SE practices with the employment of LLMs.
Code understanding. Code Understanding refers to the process of deeply comprehending and analyzing source code. It involves gaining insights into the logic, structure, functionality, and dependencies of the code (Shen et al., 2022), as well as understanding the programming languages, frameworks, and libraries used. LLMs can assist in code understanding by leveraging their powerful natural language processing capabilities to interpret code-related text, such as comments and documentation (Wang et al., 2023c; Kanade et al., 2020). They aid developers in grasping code functionality, identifying dependencies, and generating relevant code documentation (Shen et al., 2022; Ma et al., 2023a). Through their ability to comprehend both code and natural language, LLMs enhance the efficiency and accuracy of code understanding, empowering developers to maintain, optimize, and integrate code effectively (Kanade et al., 2020).
@@
Program synthesis. Program synthesis is the automated process of generating code that satisfies a given specification or set of constraints, emphasizing the derivation of functional properties of the code (Chen et al., 2017, 2021a; Manna and Waldinger, 1980; Srivastava et al., 2010; Parisotto et al., 2016). It differs from code generation, which primarily translates higher-level representations into target code without necessarily deriving its functionality from scratch (Siddiq et al., 2023a; Zhang et al., 2023e; Zheng et al., 2023b). Several studies have demonstrated that LLMs can be used for program synthesis tasks. LLMs have a significant impact on program synthesis due to their advanced language understanding and generation capabilities. LLMs can effectively interpret natural language descriptions, code comments, and requirements, and then generate corresponding code snippets that fulfill the given specifications. This helps developers in rapidly prototyping code and automating repetitive coding tasks (Kuznia et al., 2022; Gandhi et al., 2023). When applied to program synthesis, LLMs enhance productivity and reduce the burden on developers by automating the code-writing process based on high-level input (Jain et al., 2022). Their ability to understand the nuances of both natural language and programming languages makes them valuable tools in advancing the field of SE and streamlining the development lifecycle.
@@
API recommendation. Several methods have been proposed to automate API (Application Programming Interface) recommendations (Gu et al., 2016; Huang et al., 2018; Liu et al., 2018; Nguyen et al., 2016), falling into two orthogonal approaches: information retrieval-based (IR-based) and neural-based. In this context, our focus is on the latter. Wei et al. (Wei et al., 2022a) introduced CLEAR, an API recommendation method that employs the BERT sentence embedding model to represent queries, capturing continuous semantic information. Through contrast training, CLEAR enables BERT to learn precise semantic representations of queries, independent of their lexical content. Recently, Zhang et al. (Zhang et al., 2023d) developed ToolCoder, which combines API search tools with existing models to aid in code generation and API selection. This approach involves an automated data annotation method using ChatGPT, adding tool usage information to the source code data, followed by fine-tuning the code generation model. During inference, an API search tool is integrated into the generation process, allowing the model to automatically utilize the tool for suggestions when selecting APIs.
@@
API synthesis. The automated generation of application programming interface calls, known as API synthesis, plays a crucial role in bridging human intent with machine execution. In recent studies, Wang et al. (Wang et al., 2023b) and Patil et al. (Patil et al., 2023) have both explored the potential of LLMs in this realm. Utilizing models like GPT-4 and LLaMA-based architectures, these researchers showcase the prowess of LLMs in generating accurate API calls and adapting to real-time documentation changes, effectively addressing challenges like hallucination and inaccurate input arguments. The integration of LLMs in API synthesis signifies a paradigm shift, promising enhanced accuracy, adaptability, and reliability in code generation. As illuminated by these studies, the future of API synthesis may be deeply anchored in advanced machine learning, heralding new research avenues and refinements for more seamless human-machine interactions.

### v6

@@ 5.3. How are evaluation metrics utilized to assess the performance of LLM4SE tasks?
In generation tasks, metrics like BLEU, along with its variants BLEU-4 and BLEU-DC (Ahmed et al., 2024; Al-Kaswan et al., 2023; Arakelyan et al., 2023; Chen et al., 2022; Ciniselli et al., 2021), and Pass@k (Bui et al., 2023; Cassano et al., 2023; Chen et al., 2023c, 2021b; Dibia et al., 2022; Döderlein et al., 2022) are the most commonly used, appearing in 62 and 54 studies, respectively. For instance, Wang et al. (Wang et al., 2023e) employed BLEU to evaluate a code-to-code translation model. Pass@k is used in the research by Jiang et al. (Jiang et al., 2023a) to assess code generation models, measuring the proportion of generated code snippets that match the reference solutions. Additionally, ROUGE/ROUGE-L (Ahmed et al., 2024; Al-Kaswan et al., 2023; Gao et al., 2023b; Geng et al., 2024; Li et al., 2022f; Mastropaolo et al., 2021b, 2022b; Niu et al., 2022; Zan et al., 2023a; Li et al., 2023b), METEOR (Ahmed et al., 2024; Al-Kaswan et al., 2023; Chen et al., 2022; Gao et al., 2023b; Niu et al., 2022; Geng et al., 2024), EM (Exact Match) (Al-Kaswan et al., 2023; Gao et al., 2023b; Gupta et al., 2023; Murali et al., 2023; Wang et al., 2023e; Weyssow et al., 2023b; Ye et al., 2023; Zhang et al., 2023j), and ES (Edit Similarity) (Liu et al., 2023l) are used in specific studies to evaluate the quality and accuracy of generated code or natural language code descriptions.
RQ3 - Summary
(1) We discovered a range of tuning techniques gradually becoming widely adopted in the LLM4SE domain. Among these, Parameter Efficient Fine-Tuning (PEFT) techniques, including Low-Rank Adaptation (LoRA), prompt tuning, prefix tuning, and adapter tuning, are gaining prominence for optimizing LLMs while minimizing computational complexity. (2) We identified a diverse set of eight prompting techniques, including few-shot prompting, zero-shot prompting, Chain-of-Thought (CoT), Automatic Prompt Engineer (APE), Chain of Code (CoC), Automatic Chain-of-Thought (Auto-CoT), Modular-of-Thought (MoT), and Structured Chain-of-Thought (SCoT), applied in the LLM4SE domain to enhance model performance. These techniques leverage task-specific instructions, known as prompts, to guide LLMs without modifying core model parameters, providing a promising avenue for improving LLM capabilities in software engineering tasks. (3) We summarized the most widely used evaluation metrics according to four problem types, i.e., regression, classification, recommendation, and generation. Nineteen different evaluation metrics appeared in the generation task, while nine metrics were used for the classification task.
@@ 6.1. What are the distributions SE activities and problem types addressed to date with LLM4SE?
The highest number of studies is observed in the software development domain, constituting approximately 56.65% of the total research volume. This underscores the primary focus to date on utilizing LLMs to enhance coding and development processes. Software maintenance tasks account for about 22.71% of the research share, highlighting the significance of LLMs in aiding software updates and improvements. The software quality assurance domain holds approximately 15.14% of the research proportion, indicating a growing interest in automating testing procedures. In contrast, requirements engineering and software design activities represent approximately 3.9% and 0.92% of the research share, respectively, suggesting relatively limited exploration so far in these areas. The software management domain has the least research representation, accounting for a tiny 0.69% proportion. This distribution underscores the vital focus on development and maintenance tasks while also indicating potential avenues for further research in testing, design, and management domains.
In our collection of LLM studies for SE tasks, we’ve classified them based on the type of problems they address (shown in Fig.10 (b)). The distribution reveals that the majority of studies, about 70.97%, center around generation tasks, showcasing the significance of LLMs in producing code or text. Following this, around 21.61% of studies fall under classification tasks, indicating the relevance of LLMs in categorizing software elements. Additionally, roughly 6.77% of studies are related to recommendation tasks, demonstrating the utility of LLMs in suggesting solutions. Lastly, a smaller portion, around 0.65%, is allocated to regression tasks, reflecting the limited exploration of LLMs for predictive modeling. This distribution underscores the broad applicability of LLMs across different SE challenges, with a notable emphasis on code generation and classification tasks.
@@
Requirements Anaphoric ambiguity treatment (4) Requirements classification (4) 17
engineering Requirement analysis and evaluation (2) Specification generation (2)
Coreference detection (1) Requirements elicitation (1)
Specification formalization (1) Traceability automation (1)
Use cases generation (1)
Software design GUI retrieval (1) Rapid prototyping (1) 4
Software specification synthesis (1) System design (1)
Software development Code generation (118) Code completion (22) 247
Code summarization (21) Code search (12)
Code translation (12) Code understanding (8)
API inference (5) Program synthesis (6)
API recommendation (5) Code editing (5)
Code representation (3) Code comment generation (2)
Method name generation (2) Code recommendation (2)
Agile story point estimation (1) API documentation augment (1)
API documentation smells (1) API entity and relation extraction (1)
Data analysis (1) Fuzz driver generation (1)
Control flow graph generation (1) Identifier normalization (1)
Instruction generation (1) Type inference (1)
Others (14)
Software quality Vulnerability detection (18) Test generation (17) 66
assurance Bug localization (5) Verification (5)
Testing automation (4) Fault localization (3)
Defect detection (2) GUI testing (2)
Static analysis (2) Binary taint analysis (1)
Compiler fuzzing (1) Decompilation (1)
Invariant prediction (1) Malicious code localization (1)
Mobile app crash detection (1) Resource leak detection (1)
Test prediction (1)
Software maintenance Program repair (35) Code clone detection (8) 99
Code review (7) Debugging (4)
Bug reproduction (3) Review/commit/code classification (3)
Duplicate bug report detection (3) Logging (3)
Log parsing (3) Code revision (2)
Sentiment analysis (3) Vulnerability repair (2)
API misuses repair (1) Bug prediction (1)
Bug triage (1) Code coverage prediction (1)
Code review explained (1) Code-Review defects repair (1)
Crash bug repair (1) Dockerfile Repair (1)
Incivility detection (1) Patch correctness prediction (1)
Patch detection (1) Program merge conflicts repair (1)
Rename Refactoring (1) Tag recommendation (1)
Technical debt payback (1) Traceability recovery (1)
Web test repair (1) Type error repair (1)
Others (5)
Software management Effort estimation (2) Software tool configuration (1) 3
  • *See Appendix E for the full table including references.
@@ 6.2. How are LLMs used in requirements engineering?
Anaphoric ambiguity treatment. Ambiguity in software requirements arises when a single reader can interpret a natural language (NL) requirement in multiple ways, or when different readers have varying understandings of the same requirement. Unclear and ambiguous NL software requirements can lead to suboptimal software artifacts during later development stages. Moharil et al. (Moharil and Sharma, 2023) and Ezzini et al. (Ezzini et al., 2022) have empirically demonstrated the significant role of LLMs such as BERT and SpanBERT in effectively addressing anaphoric ambiguity. Sridhara et al. (Sridhara et al., 2023) revealed that ChatGPT excels in addressing anaphoric ambiguity in software requirements. Through researchers’ analysis of ten English requirement specifications (Ezzini et al., 2022) containing anaphora-related challenges, ChatGPT consistently demonstrated its remarkable capability to accurately identify antecedents. This empirical evidence emphasizes the valuable role ChatGPT can play in enhancing the clarity and precision of software requirements, ultimately contributing to more effective software development processes by reducing interpretational uncertainties.
@@
Others. In addition to the four requirement engineering tasks detailed above, LLMs can also be applied to requirement analysis and evaluation (Poudel et al., 2023; Ronanki et al., 2023), specification generation (Ma et al., 2024a; Xie et al., 2023b), requirements elicitation (White et al., 2023b), specification formalization (Endres et al., 2023), and use case generation (Zhang et al., 2024d).
@@ 6.4. How are LLMs used in software development?
Code generation. Code generation has long been a task of interest: there is extensive work on program synthesis using symbolic and neural-semiotic approaches (Alur et al., 2013; Wu et al., 2023b). Recently, LLMs trained for text generation have demonstrated the ability to complete programs (Brown et al., 2020; Black et al., 2022). Since 2020, several code generation models have been trained or fine-tuned on programming language text (Nijkamp et al., 2022b; Chen et al., 2021b; Fried et al., 2022; Xu et al., 2022; Feng et al., 2020; Clement et al., 2020). Unlike traditional program synthesis techniques, neurolinguistic models can be conditioned on natural language (e.g., code annotations) as well as generate programming language text. Researchers have experimentally demonstrated that LLMs like GPT-4 (Bareiß et al., 2022; Liu et al., 2023k; Jiang et al., 2023c; Gilbert et al., 2023), GPT-2/GPT-3/GPT-3.5 (Azaria et al., 2023; Yetiştiren et al., 2023; Ke et al., 2023; Liu et al., 2023k; Nascimento et al., 2023; Li et al., 2023c; Wang et al., 2023h; Liu et al., 2023a; Dong et al., 2023b), BERT series (Zeng et al., 2022; Lai et al., 2023), Codex (Dibia et al., 2022; Bareiß et al., 2022; Yu et al., 2023a; Chen et al., 2021b; Gupta et al., 2023; Madaan et al., 2022; Kuznia et al., 2022), CodeGen (Dibia et al., 2022; Jones and Steinhardt, 2022; Zan et al., 2022a), InCoder (Murali et al., 2023; Kou et al., 2023b; Liu et al., 2023k; Wang et al., 2022b), Copilot (Wu et al., 2023b) and CodeGeeX (Zheng et al., 2023c), play a key role in code generation. By pre-training on large-scale text data, these models learn rich linguistic knowledge and semantic representations that enable them to understand the meaning and structure of natural language. LLMs can automate code generation by converting natural language descriptions into code (Jiang et al., 2023a). These models generate program code from natural language descriptions, enhancing code-writing efficiency and accuracy. They show excellent performance in code completion, automatic code generation, and conversion of natural language annotations to code, providing software developers with powerful auxiliary tools and promoting further automation and intelligence in the code writing and development process.
@@
Within the domain of LLMs applied to software development tasks, studies centered on code generation distinctly dominate the academic landscape. As reflected in Table 11, the GPT series, particularly GPT-4, emerged as a key focus, with many more studies using them in the realm of code generation (Dong et al., 2023b; Du et al., 2023b; Li et al., 2023c; Liu et al., 2023k). Analyzing these studies, several noteworthy findings surface:
@@
  • Programming thinking in LLMs. Techniques that evoke “programming thinking” within LLMs, such as the TIP (i.e., Thinking in Programming) (Li et al., 2023c) methodology, have shown promising strides. By guiding LLMs to first craft a high-level code sketch before delving into detailed implementations, the synthesized code exhibits higher accuracy and robustness.
@@
  • Class-level vs. Method-level generation. LLMs, while adept at method-level code generation, present varied performance metrics when tasked with class-level generation (Du et al., 2023b). This divergence underscores the evolving nature of challenges as the granularity of code synthesis shifts.
@@
  • Expanding LLM capabilities. The next frontier in this discipline seems to lie in harmoniously integrating LLMs with established SE tools and practices. The emergence of frameworks like EvalPlus (Dong et al., 2023b) indicates a trend towards enhancing the evaluation and accuracy of LLM-generated code, possibly ushering in an era where human developers and LLMs collaboratively craft software solutions.
@@
GPT-3.5
Codex, CodeGen, CodeGeeX, LLaMA, InCoder, PyCodeGPT, CodeParrot, GPT-2
HumanEval, MBPP, MBCPP
Pass@k May 11, 2023 (Li et al., 2023c)
GPT-4
PaLM Coder, Codex, CodeGen-Mono, Incoder, CodeGeeX, AlphaCode
HumanEval, HumanEval-ET, MBPP, MBPP-ET
Pass@k May 24, 2023 (Dong et al., 2023b)
GPT-4
GPT-3.5, StarCoder, CodeGen, CodeGen2, Vicuna, SantaCoder, Incoder, GPT-J, GPT-Neo, PolyCoder, StableLM
HumanEval, HumanEval+, HumanEval-mini
Pass@k Jun 12, 2023 (Liu et al., 2023k)
GPT-4
GPT-3.5, WizardCoder, Instruct-StarCoder, SantaCoder, Instruct-CodeGen, CodeGeeX, InCoder, Vicuna, ChatGLM, PolyCoder
ClassEval, HumanEval
Pass@k Aug 3, 2023 (Du et al., 2023b)
@@
With the support of LLMs, code completion achieves significant improvements in efficiency and accuracy. Developers can save time by avoiding manual input of lengthy code and reducing the risk of code errors. LLMs also learn from extensive code repositories, acquiring knowledge and best practices to offer more intelligent and precise suggestions, aiding developers in better understanding and utilizing code (Ciniselli et al., 2021). Additionally, these models can provide personalized code recommendations based on developers’ coding styles and preferences, further enhancing the effectiveness and user experience of code completion (Liu et al., 2023l).
Code summarization. Code summarization is a task that attempts to understand the code and automatically generate descriptions directly from the source code. It can also be viewed as an extended form of documentation. Successful code summarization not only facilitates the maintenance of source code (Iyer et al., 2016; Nguyen and Nguyen, 2017) but can also be used to improve the performance of code search using natural language queries (Nie et al., 2016; Yang et al., 2016) and code classification (Nguyen and Nguyen, 2017). LLMs play a significant role in code summarization by analyzing code structures and contexts to generate informative natural language summaries. Specifically, LLMs such as Codex (Gao et al., 2023b; Arakelyan et al., 2023; Ahmed et al., 2023), CodeBERT (Chen et al., 2022; Gu et al., 2022; Gao et al., 2023b), and T5 (Mastropaolo et al., 2021b, 2022b) comprehend the functionality and logic of the code, producing easily understandable human language descriptions. For example, Arakelyan et al. (Arakelyan et al., 2023) rigorously evaluate the efficacy of CodeT5 and Codex across code generation and summarization tasks, shedding light on their performance under distribution shifts. It unveils practical adaptation techniques, underscoring Codex’s commendable performance. Additionally, the study demonstrates that while adapted models exhibit proficiency in code generation, their generality can present trade-offs in the context of code summarization. As a result, code summarization with the support of LLMs enhances code readability, improves software documentation quality, and accelerates code comprehension and collaboration among developers. This advanced approach to code summarization demonstrates great potential for automating and streamlining various aspects of software development in modern SE practices with the employment of LLMs.
@@
Code understanding. In contrast to code summarization, which focuses on automatically generating human-readable descriptions from source code, code understanding involves a deep analysis of source code to comprehend its logic, structure, functionality, and dependencies, as well as understanding the programming languages, frameworks, and libraries used (Shen et al., 2022). LLMs can assist in code understanding by leveraging their powerful natural language processing capabilities to interpret code-related text, such as comments and documentation (Wang et al., 2023e; Kanade et al., 2020). They aid developers in grasping code functionality, identifying dependencies, and generating relevant code documentation (Shen et al., 2022; Ma et al., 2023a). Through their ability to comprehend both code and natural language, LLMs enhance the efficiency and accuracy of code understanding, empowering developers to maintain, optimize, and integrate code effectively (Kanade et al., 2020).
@@
Program synthesis. Program synthesis is the automated process of generating code that satisfies a given specification or set of constraints, emphasizing the derivation of functional properties of the code (Chen et al., 2017, 2021a; Manna and Waldinger, 1980; Srivastava et al., 2010; Parisotto et al., 2016). It differs from code generation, which primarily translates higher-level representations into target code without necessarily deriving its functionality from scratch (Siddiq et al., 2023a; Zhang et al., 2023l; Zheng et al., 2023c). Several studies have demonstrated that LLMs can be used for program synthesis tasks. LLMs have a significant impact on program synthesis due to their advanced language understanding and generation capabilities. LLMs can effectively interpret natural language descriptions, code comments, and requirements, and then generate corresponding code snippets that fulfill the given specifications. This helps developers rapidly prototype code and automate repetitive coding tasks (Kuznia et al., 2022; Gandhi et al., 2023). When applied to program synthesis, LLMs enhance productivity and reduce the burden on developers by automating the code-writing process based on high-level input (Jain et al., 2022). Their ability to understand the nuances of both natural language and programming languages makes them valuable tools in advancing the field of SE and streamlining the development lifecycle.
@@
API recommendation. Several methods have been proposed to automate API (Application Programming Interface) recommendations (Gu et al., 2016; Huang et al., 2018; Liu et al., 2018; Nguyen et al., 2016), falling into two orthogonal approaches: information retrieval-based (IR-based) and neural-based. In this context, our focus is on the latter. Wei et al. (Wei et al., 2022a) introduced CLEAR, an API recommendation method that employs the BERT sentence embedding model to represent queries, capturing continuous semantic information. Through contrast training, CLEAR enables BERT to learn precise semantic representations of queries, independent of their lexical content. Recently, Zhang et al. (Zhang et al., 2023k) developed ToolCoder, which combines API search tools with existing models to aid in code generation and API selection. This approach involves an automated data annotation method using ChatGPT, adding tool usage information to the source code data, followed by fine-tuning the code generation model. During inference, an API search tool is integrated into the generation process, allowing the model to utilize the tool for suggestions when selecting APIs automatically.
@@
API inference. The automated generation of application programming interface calls, known as API synthesis, plays a crucial role in bridging human intent with machine execution. In recent studies, Wang et al. (Wang et al., 2023d) and Patil et al. (Patil et al., 2023) have both explored the potential of LLMs in this realm. Utilizing models like GPT-4 and LLaMA-based architectures, these researchers showcase the prowess of LLMs in generating accurate API calls and adapting to real-time documentation changes, effectively addressing challenges like hallucination and inaccurate input arguments. The integration of LLMs in API synthesis signifies a paradigm shift, promising enhanced accuracy, adaptability, and reliability in code generation. As illuminated by these studies, the future of API synthesis may be deeply anchored in advanced machine learning, heralding new research avenues and refinements for more seamless human-machine interactions.